Assume the operating income is 500,000, interest rate is 10%, outstanding debt is 2,000,000 and WACC=12.5%

Required;

1.The total value of the firm

2.The market value of equity

3.Cost of equity
Answer #1
  • Operating Income = 500,000
  • Less: Interest = 200,000 [being 10% of 2,000,000]
  • Net Profit = 300,000 [Tax effects are ignored. If tax rates are given, it shall be deducted too.]

1. Total value of the firm = 300,000 / 12.5% = 2,400,000

2. Market value of equity = 2,400,000 - 2,000,000 = 400,000

3. Cost of equity = [12.5% - (10% x 2 / 2.4)] x 2.4 / 0.4 = 25%

Hope this helps.
Please leave a thumbs up.
Thank you!